Rose's garden is in the shape of a trapezoid. If the height of the trapezoid is 16m, one base is 20m, and the area is 224m(squared by 2), find the length of the other base.
Found 2 solutions by stanbon, sarah_adam:
Answer by stanbon(48569)

Area= (1/2)[b1+b2)h
224 = 8(b1 +20)
28 =b+20
base = 8 ft
Cheers,
stan H.

Answer by sarah_adam(201)

If a and b are the lengths of the two parallel bases of a trapezoid, and h is its height, the area of the trapezoid is
1/2 × h × (a + b) .
here h = 16 m
length of one base a = 20 m
area = 224 m^2
224 = 1/2 * 16* (20 +b)
224 = 8*(20+b)
224/8 = 20+b
28 = 20 + b
b = 28 - 20 = 8m
Length of another base is 8m